OpenStudy (anonymous):

Problem Set 1, section 1A, problem 6b: Can someone please explain how it is determined that Acos(y)=1 and Asin(y)=-1? I didn't see any of this covered in any of the independent study lecture videos, so I'm assuming it's review material (if it is, I haven't seen it in a long time). Thanks.

OpenStudy (anonymous):

can you state the full problem?

OpenStudy (anonymous):

Express the following in the form A*sin(x+c): sinx-cosx

OpenStudy (anonymous):

From the Sum Identity http://www.mathportal.org/algebra/trigonometry/trigonometric-formula.php we know that sin (x + y) = sin(x) cos(y) + cos(x) sin(y) we can get a similar form as the required form (substitute c = y and multiply both sides with A A*sin (x + c) = A*sin(x) cos(c) + A*cos(x) sin(c) we need the form sin(x)-cos(x) A*sin (x + c) = A*sin(x) cos(c) + A*cos(x) sin(c) = sin(x)-cos(x) Hence  as you can see from the drawing wee need to solve the system of equations A*cos(c) = 1 A*sin(c) = -1 (which explains your original question!) This is equaivalent to cos(c) = 1/A sin(c) = -1/A then from the first we can see that A = 1/cos(c) and substitute in in the second then we get: sin(c)/cos(c) = -1 or sin(c) =- cos(c), and this is at c = -pi/4 or 3*pi/4 now cos(-pi/4) = 1/sqrt(2) so A = sqrt(2), which has to check with sin((-pi/4) =-/sqrt(2) so A = sqrt(2) indeed So we have the first solution: sqrt(2) * sin(x-pi/4) = sin(x)-cos(x) or cos(3pi/4) = -1/sqrt(2) so A = -sqrt(2), which has to check with sin((3pi/4) =1/sqrt(2) so A =-sqrt(2) indeed So the second solution is -sqrt(2) * sin(x+3pi/4) = sin(x)-cos(x) This can be shown in a graph

OpenStudy (anonymous):

@ERMiller

OpenStudy (anonymous):

Thanks for your help. I still don't see why A*cos(c) = 1 and A*sin(c) = -1. It seems like you arrived at that first and went from there. I understand the process once we determine A*cos(c) = 1 and A*sin(c) = -1, I just don't get how that is determined.

OpenStudy (anonymous):

This is the case because of the general property sin (x + y) = sin(x) cos(y) + cos(x) sin(y) Do you know how that property? (Have you ever seen how that was proven?)

OpenStudy (anonymous):

Anyway, if you use that 'law', you can see that if sin(x) - cos(x) is of the form A*(sin(x) cos(y) + cos(x) sin(y)), it means you is equal to A*sin(x+y) sin(x) - cos(x) is only of that form if A*cos(y) = 1 and A*sin(y) = 1 if you backsubstitute it into A*(sin(x) cos(y) + cos(x) sin(y)) = A*cos(y) sin(x) +A*sin(y)cos(x) we get the needed sin(x)-cos(x) For A*cos(y) = 1 and A*sin(y) = 1 to be true then y = -pi/4 and A = sqrt(2) or y = 3*pi/4 and A = -sqrt(2) From the graph you can see that that is the right solution (whether you use y or c makes no difference)

OpenStudy (anonymous):

Here is a nice proof of the identities like sin (x + y) = sin(x) cos(y) + cos(x) sin(y) http://www.youtube.com/watch?v=zcEMKv5yIYs

OpenStudy (anonymous):

Oh, ok - I had never seen it proven despite being aware of the property. Thanks so much for taking all that time to explain it. I really appreciate it.

OpenStudy (anonymous):

It's a good repetition for me too, I hadn't looked at the proof since secondary school andymore. These and more trigonometry identities are used very often. It often helps to just think "Is there any one of those I can use?"

OpenStudy (anonymous):

A*cos(y) = 1 and A*sin(y) = 1 just fits the bill